# A Digital twin for Diesel Engines: Operator-infused PINNs with Transfer Learning for Engine Health Monitoring

### v1

@@ 2.4 A hybrid model using DeepONet and PINN for the engine problem
The individual physics losses are represented as:
$\mathcal{L}_{p_{im}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(p_{im}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(\dfrac{dp_{im}^{(k)}}{dt}-\dfrac{R_{a}T_{im}}{V_{im}}\left(W_{c}^{(k)}+W_{egr}^{(k)}-W_{ei}^{(k)}\right)\right)^{2}$ (21a)
$\mathcal{L}_{p_{em}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(p_{em}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(\dfrac{dp_{em}^{(k)}}{dt}-\dfrac{R_{e}T_{em}^{(k)}}{V_{em}}\left(W_{eo}^{(k)}-W_{t}^{(k)}-W_{egr}^{(k)}\right)\right)^{2}$ (21b)
$\mathcal{L}_{\omega_{t}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(\omega_{t}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(\dfrac{d\omega_{t}^{(k)}}{dt}-\dfrac{P_{t}^{(k)}\eta_{m}-P_{c}^{(k)}}{J_{t}\omega_{t}^{(k)}}\right)^{2}$ (21c)
$\mathcal{L}_{x_{r}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(x_{r}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(x_{r}^{(k)}-\dfrac{(\Pi_{e}^{(k)})^{1/\gamma_{a}}(x_{p}^{(k)})^{-1/\gamma_{a}}}{r_{c}x_{v}^{(k)}}\right)^{2}$ (21d)
$\mathcal{L}_{T_{1}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(T_{1}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(T_{1}^{(k)}-\left(x_{r}^{(k)}T_{e}^{(k)}+(1-x_{r}^{(k)})T_{im}\right)\right)^{2}$ (21e)
$\mathcal{L}^{data}_{\phi}(\bm{\lambda}_{\phi})=$ $\dfrac{1}{n_{\phi}}\sum_{j=1}^{n_{\phi}}\left[\left(\phi_{data}^{(j)}-\hat{\phi}_{data}^{(j)}\right)\lambda_{\phi}^{(j)}\right]^{2},\quad\quad\phi=p_{im},p_{em},\omega_{t},W_{egr}$ (22a)
$\mathcal{L}^{data}_{T_{em}}=$ $\dfrac{1}{n_{T_{em}}}\sum_{j=1}^{n_{T_{em}}}\left[\left(T_{em_{data}}^{(j)}-\hat{T}_{em_{data}}^{(j)}\right)\right]^{2}$ (22b)
The initial condition losses are given as:
$\mathcal{L}^{ini}_{\phi}=\dfrac{1}{1}\sum_{j=1}^{1}\left(\phi_{0}^{(j)}-\hat{\phi}_{0}^{(j)}\right)^{2},\;\;\;\;\;\;\;\;\phi=p_{im},p_{em},\omega_{t},x_{r},T_{1}$ (23)
The parameters of the neural networks for the hybrid PINN model (ref. Table 4), along with unknown parameters (ref. Table 1) and the self-adaptive weights ($\lambda$), are obtained by optimizing the loss function (20) using the Adam [31] optimizer. The optimization process occurs in two phases. In the first phase,we optimize the PINN network parameters, unknown parameters, and self-adaptive weights over a certain number of epochs. At the end of this phase, we freeze the values of the self-adaptive weights. Thereafter, we optimize only the parameters of the hybrid PINN model and the unknown parameters using Adam, keeping the self-adaptive weight constant with the values obtained at the end of the first phase. As discussed earlier, the parameters of the pre-trained neural networks for the empirical formula and the DeepONets are kept frozen during the hybrid PINN optimization process (both phases).
@@ 2.5 Multi-stage Transfer learning
In the preceding sections, we have outlined a method for predicting unknown parameters and gas flow dynamics of a diesel engine. In this method, the parameters of the networks, unknown parameters and the self-adaptive weights are initialized randomly and trained to obtain the optimal parameters. Due to the stochastic initialization of the network parameters, unknown parameters and self-adaptive weights, significant computation time is required to converge to an accurate estimate. Furthermore, any changes in the control input vector $u=\{u_{egr},u_{vgt},u_{\delta}\}$, engine speed $n_{e}$ or ambient conditions necessitates re-training the entire PINN model.
@@
To mitigate the challenge related to computational time, we implement a multi-stage transfer learning to train the hybrid PINN model separately for each one-minute time segment as shown in figure 3. Transfer learning [32, 33, 34] utilizes knowledge gained from training a model on one dataset (typically referred to as source data) to enable knowledge transfer to train on a different but similar dataset (known as target data). We demonstrate our methodology using a five-minute long data signal divided into five one-minute segments. First, the complete hybrid PINN model is trained with the given control vector $\{u_{\delta},u_{egr},u_{vgt}$}, and engine speed $n_{e}$ for time segment $0-60$s. The weights and biases of the network, denoted by $\bm{\theta}$, are initialized randomly. The unknown parameters for the engine, $\bm{\Lambda}:=\{A_{egrmax},A_{vgtmax},\eta_{sc},h_{tot}\},\bm{\Lambda}\in\mathbb{R}^{+}$ are initialized with a random positive value (discussed in detail in section 4). Similarly, the self-adaptive weights $\bm{\lambda}$ are also initialized randomly at the beginning of the training. Thereafter, the PINN network is trained as per the process outlined in section 2.4, and the network parameters and unknown parameters $\bm{\Lambda}$ are saved. In the transfer learning stage, these saved parameter values are used for initialization of network parameter (weights and biases) and unknown parameters, i.e. instead of random initialization, we import these values from the trained model in $0-60$ sec. In Figure 3, we illustrate different stages of the proposed transfer learning scheme, and these stages include:
  1. Stage I:
    Transfer weights and biases ($\bm{\theta}$) and parameter estimated ($\bm{\Lambda}$) from the trained model on segment 0-60 sec to the new training in new time-segment. Initialize self-adaptive weights ($\bm{\lambda}$) randomly. Train complete PINN model and update $\bm{\theta},\bm{\Lambda}$, $\bm{\lambda}$} upto $N_{1}$ epochs.
  2. Stage II:
    Freeze all network weights ($\bm{\theta_{l-1}}$) except the last layer ($\bm{\theta_{l}}$). Update $\bm{\theta_{l}},\bm{\Lambda}$, $\bm{\lambda}$} upto $N_{2}$ epochs.
  3. Stage III:
    Freeze $\bm{\theta_{l-1}}$ and $\bm{\lambda}$, update $\bm{\theta_{l}},\bm{\Lambda}$} upto $N_{3}$ epochs.
@@ 3. Stage III:
Fig. 3: Schematic for Multi-stage transfer learning showing the network training protocol for different segments of time. In this approach, the hybrid PINN model is trained for the first segment of time (0-60 sec). Thereafter, for the subsequent time segments (60-300s in each one-minute segment), we transfer the network parameters ($\bm{\theta}$) and engine parameter estimated ($\bm{\Lambda}$) from the hybrid PINN model trained on the first time segment. Each segment uses a different control vector $\{u_{\delta},u_{egr},u_{vgt}$}, and engine speed $n_{e}$. The adaptive weights ($\bm{\lambda}$) are initialized at the start of training at epoch=0. The network weights and biases of all the layers ($\bm{\theta}$), self-adaptive weights ($\bm{\lambda}$), and unknown parameters ($\bm{\Lambda}$) are updated in Stage I training phase. In Stage II, only weights and biases of the last layer of the PINN model ($\bm{\theta_{l}}$) are trained while freezing all the hidden layers ($\bm{\theta_{l-1}}$). Additionally, $\bm{\lambda}$ and $\bm{\Lambda}$ continue to be updated in this phase. In Stage III, we continue to update $\bm{\theta_{l}}$ and $\bm{\Lambda}$ while the self-adaptive weights $\bm{\lambda}$ are fixed to the values obtained at the end of Stage II. This approach avoids the need to re-train the complete model again and helps reduce the computation time for parameter estimation. In total, for predicting across a 5-minute time window, we train five models, with the model trained for segment 1 (0-60 sec) serving as the reference model from which the remaining models acquire knowledge during the transfer learning phase.
@@ 2.6 Few-shot transfer learning
In the previous sections, first we developed a hybrid DeepONet and PINN model and proposed a multistage transfer learning for the engine inverse problem discussed in section 1.2. In the method, for the first time segment, the model needs to be trained from the beginning with random initialization of the weights and biases. Secondly, to calculate the physics loss, the derivative for the output of the neural network with respect to the input needs to be calculated at each epoch. Thus, although transfer learning improves the computation time, it is still high as per application requirements. One approach to improve computational time is by reducing (or eliminating) the differentiation operation involved in PINN training . This may be achieved by considering the derivative to be a linear combination of some base vector. For example, in function connection [35], multi-head network [36] or reservoir computing [37], we generally do not train the hidden layer and approximate the derivative of the output as a linear combination of the derivative of the hidden layer. Desai et al. [36] incorporated one-shot inference on new tasks by training the PINN network on a training task set and then freezing all except the last layer that is trained for the new task. This approach enables one to obtain the network output as a linear matrix multiplication of the output from the frozen layer with the weight matrix of the last layer. This reduces the need for automatic differentiation to evaluate the gradients of outputs with respect to inputs, which is computationally expensive. We adopt this transfer learning paradigm here, called few-shot transfer learning.
@@
Fig. 4 illustrates a schematic diagram of a few-shot transfer learning combined with pre-trained DNN and DeepONet. Here, we note that the objectives of the pre-trained DNN and DeepONet are the same as discussed in the previous section. Hybrid PINN, using a few shot transfer learning, may be divided into two phases. The first phase (Phase I) is an offline training phase with multiple heads for unknown variables. In this phase, we trained the network with multiple heads, each head approximating one minute (corresponding to a different time zone) of the same output variable. The input to the network is time $0-60$ sec, normalized between $[-1,1]$. The first phase of offline training aims to obtain good universal basis functions (hidden layers), which can be combined to obtain appropriate output in Phase II for solving the inverse problem using PINN. In the first phase, the network is trained in a data-driven setting in an offline mode. Once the network is trained, the hidden layers (universal basis function) are transferred to the hybrid PINN model to estimate the variables in Phase II.
@@
In Phase II, we transfer the hidden layer of the multi-head network trained in the first phase. We randomly initialize the last layer ($\bm{W}_{out},b_{out}$) and setup the hybrid PINN inverse problem for one minute duration independently. In this phase, only the last layer of the network and the unknown parameters of the equation and the self-adaptive weights are optimized. The estimated output is obtained as
@@
where $\bm{h}(t)$ is the output of the last hidden layer, $\bm{W}_{out}$ and $b_{out}$ are the weights and biases of the output layer. The derivative of the output w.r.t. the input variable is
@@
Since in Phase II, the hidden layers are fixed (not trained), the derivative $\dfrac{d\bm{h}(t)}{dt}$ does not change and can be kept outside the training loop. Thus, the derivatives need to be calculated only once outside the training loop. In other words, we approximate the derivative of the output as a linear combination of the derivative of the hidden layer with the same combination matrix ($\bm{W}_{out}$) as that of the output. The loss function for the few shot transfer learning is the same as that of hybrid PINN given by Eq. (20). Note here that we utilize an output transformation function and hence we have to employ the chain rule for calculating the derivative of output with respect to input. We discuss this in detail in the numerical result section 5.
Fig. 4: Schematic diagram for the few-shot transfer learning architecture with multi-head network. The left side of the plot shows Phase I of the few-shot transfer learning approach. In this phase, we train the multi-head networks using laboratory data where each head approximates one minute of response (corresponding to a different time zone). Input to this network is time $0-60$ sec, normalized between $[-1,1]$. Phase I training is a data-driven offline training. On the right side of the plot, Phase II that involves transfer learning is shown. In this phase, we transfer the hidden layer and derivative of the output of the last hidden layer from the network trained in Phase I. We randomly initialize the weights and biases of the last layer along with the unknown parameters. The output and the derivative of the output are predicted using Eqs. 24 and 25, respectively. We formulate the hybrid PINN model by combining the PINN network with DeepONet and empirical DNN. The total loss ($\mathcal{L}_{\bm{\theta}}$) is calculated as the weighted sum of the physics and data loss. The weights $\lambda_{1}$ and $\lambda_{2}$ may be constant or adaptive depending on the problem setup and solution method. The trainable parameters are the weights and bias of the output layer of the PINN network ($\bm{\theta}_{out}=\{\bm{W},b\}$), the unknown parameters of the equation, the self-adaptive (in case of self-adaptive) weights. The weight and biases of the hidden layer of the PINN network, the parameters of the pre-trained network for empirical formula, and the parameters of the DeepONets remain fixed during the optimization process.
@@ 2.7 Data generation
As discussed earlier, we utilize two types of networks: (1) pre-trained DNN for the empirical formula and DeepONets (pre-trained) for the independent states ($\tilde{u}_{egr1}$, $\tilde{u}_{egr2}$ and $\tilde{u}_{vgt}$), (2) PINN networks used for online learning. Thus, two distinct types of data is required. We label the data used for training the empirical formulae networks and DeepONets for EGR and VGT states as laboratory data. The data used for solving the inverse problem through PINNs is labeled as field data, which is used to simulate a real testing scenario. Our prior study [20] provides an in-depth discussion on the laboratory data required for training each DNN for empirical formulae and other necessary details. In the absence of laboratory and/or field data, we use Simulink [24] to generate both datasets under different conditions as presented in Table 1.
In the present study, we consider four sets of input data (control vector and speed) to generate lab and field data as shown in Table 5. The simulated data are generated with $dt=0.2$ sec for training and testing of DNN for empirical formulae and DeepONet except training of $\eta_{tm}$. In the case of $\eta_{tm}$, these are generated at $dt=0.025$ sec. We trained the DNN for $f_{egr}(\mathcal{N}_{2}^{P}(:,\bm{\theta}))$ only with data from Case-I shown in table 5. The data for training of multi-head DNN for few-shot transfer learning are discussed later in the numerical results section. For further details regarding data generation, we refer readers to our previous study [20]. We consider PINN for each 60 sec. time segment independently in our computational study. The initial conditions for all the state variables and $x_{r}$ and $T_{1}$ are known for each time segment considered in the numerical study.
@@
(K) (Pa) $dt$
Case-I $T-65$ $P-0.100$ Set-I 1 sec Training pre-trained DNN for empirical formulae
Case-II $T-65$ $P+0.211$ Set-I 1 sec
Case-III $T-28$ $P-0.100$ Set-I 1 sec
@@
Case-I $T-65$ $P-0.100$ Set-I 1 sec Training $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$
@@
Case-I $T-65$ $P-0.100$ Set-I 1 sec Training $\tilde{u}_{vgt}$
@@
3 Results: Pre-trained networks and DeepONets
To enhance the organization and readability of our numerical results, we divide the discussion into three sections. First, we discuss the results of pre-trained DNNs used for estimating empirical relationships and the DeepONets for the EGR and VGT states. These two networks are integrated into the hybrid PINN framework. In section 4, we present the results pertaining to the application of PINNs and the multi-step transfer learning approach. In Section 5, we discuss numerical results for the few-shot transfer learning approach.
3.1 Pre-trained networks: Neural network for empirical formulae
The empirical formulae of the engine model are approximated using surrogate neural networks with dropout as discussed in section 2.2. We apply a 2% dropout for the hidden layers of these networks. Notably, dropout is employed not only during the training of these networks but also during the PINN optimization process for the inverse problem and prediction. This helps us to analyze the potential variation in PINN estimates expected due to systematic (e.g. manufacturing) variability for real systems. We also added a 3% Gaussian noise to the output state variables during training to replicate the expected noise in laboratory measurement.
3.2 DeepONet: Approximation of EGR states ($\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$) and VGT states ($\tilde{u}_{vgt}$) using DeepONets
[…]
The loss function for individual output variables is the mean square error between the true value ($\tilde{u}_{1}$) and the predicted one ($\tilde{u}_{1}(\bm{\theta})$). The total loss is the weighted sum of losses for both outputs. For EGR actuator dynamics, we use weights $a_{1}=1$ and $a_{2}=100$ (ref. Fig. 5). The DeepONets are trained using the Adam optimizer, with $3\times 10^{5}$ epochs for DeepONet $\mathcal{D}_{1}$ and $4\times 10^{5}$ epochs for DeepONet $\mathcal{D}_{2}$. The input and output variables are scaled and are as shown in Table 6 along with including the relative $L_{2}$ error for prediction. The mean and standard deviation of the error are calculated from six independent training sessions and show good accuracy. To further validate the model’s robustness, we added a $3\%$ Gaussian noise to the labelled output data and trained the networks with this noisy data. The relative $L_{2}$ error for these cases is also shown in Table 6, and we note only a marginal change in error, demonstrating the model’s robustness to noise in training data.
Fig. 5: Schematic diagram of DeepONet for prediction of EGR states showing branches and trunk. The trunk network takes time as its input, at which the value of the output function needs to be evaluated. The four branches take $u_{d}$, $\tilde{u}_{1,0}$, $u_{d}$ and $\tilde{u}_{2,0}$ as input and approximate two outputs $\tilde{u}_{1}=\tilde{u}_{egr1}$ and $\tilde{u}_{1}=\tilde{u}_{egr2}$. Here, $\tilde{u}_{1,0}=\tilde{u}_{egr1,0}$ and $\tilde{u}_{2,0}=\tilde{u}_{egr2,0}$ are the initial conditions and $u_{d}=u_{degr}=u_{egr}(t-\tau_{degr})$, where $\tau_{degr}$ is the time delay coefficient. The loss function for individual output variables is the mean square error between the true value ($\tilde{u}_{1}$) and the predicted one ($\tilde{u}_{1}(\bm{\theta})$), and the total loss is the weighted sum ($a_{1}=1$, $a_{2}=100$) of loss for both outputs. In the case of prediction of VGT actuator states, there are only two branch networks, branch 1 and 2 and only one output $\tilde{u}_{1}=\tilde{u}_{vgt}$. The inputs to these branch networks are $u_{d}=u_{dvgt}=u_{egr}(t-\tau_{dvgt})$ and the initial condition $\tilde{u}_{1,0}=\tilde{u}_{vgt,0}$. Appropriate scaling of the input and output are considered for fast and better training. Details of the network sizes and input-output data structures are shown in Table C.1.
Table 6: DeepONet results for independent variables: The training and testing errors for DeepONet for EGR and VGT states. The two DeepONets are trained using clean data and with added Gaussian noise. The relative $L_{2}$ error (%) for training and testing are shown in the fourth and sixth columns, respectively. The bottom part of the table gives the $L_{2}$ error (%) in training and testing when Gaussian noise is added to the training labelled data. We also note that the error in the case of noisy data is calculated with respect to the clean data. The mean and standard deviation of the error are calculated based on six different independent runs.
DeepONet Output Training Testing
Variable Data type Rel. $L_{2}$ error (%) Data type Rel. $L_{2}$ error (%)
$\tilde{u}_{egr2}$ Clean data $0.65\pm 0.04$ Clean data $3.64\pm 0.18$
with added noise in output during training
$\tilde{u}_{egr2}$ 3% noise added $0.97\pm 0.02$ Clear data $3.48\pm 0.17$
4 Results: PINN for inverse problem and transfer learning

### v2

@@ 2.4 A hybrid model using DeepONet and PINN for the engine problem
The optimal parameters of the neural networks for the hybrid PINN model (ref. Table 4), along with unknown parameters (ref. Table 1) and the self-adaptive weights ($\lambda$), are obtained by optimizing the loss function from Eq. (18) using the Adam [41] optimizer. The optimization process is divided into two phases. In the first phase, we optimize the PINN network parameters, unknown parameters, and self-adaptive weights for a certain number of epochs. At the end of this phase, we freeze the values of the self-adaptive weights. Thereafter, we optimize only the parameters of the hybrid PINN model and the unknown parameters using Adam, keeping the self-adaptive weight constant with the values obtained at the end of the first phase. As discussed earlier, the parameters of the pre-trained neural networks for the empirical formula and the DeepONets are kept frozen during the hybrid PINN optimization process (both phases).
@@ 2.5 Multi-stage Transfer learning
In the preceding sections, we have outlined a method for predicting unknown parameters and gas flow dynamics of a diesel engine. In this method, the parameters of the networks, unknown parameters and the self-adaptive weights are initialized randomly and trained to obtain the optimal parameters. Due to the stochastic initialization of the network parameters, unknown parameters and self-adaptive weights, significant computation time is required to converge to an accurate estimate. Furthermore, any changes in the control input vector $\bm{u}=\{u_{egr},u_{vgt},u_{\delta}\}$, engine speed $n_{e}$ or ambient conditions necessitates re-training the entire PINN model.
@@
To mitigate the challenge related to computational time, we implement a multi-stage TL paradigm to train the hybrid PINN model individually for each one-minute time segment as shown in Fig. 3. Transfer learning [42, 43, 44] utilizes knowledge gained from training a model on one dataset (typically referred to as source data) to enable knowledge transfer to train on a different but similar dataset (known as target data). We demonstrate our methodology using a five-minute long data signal divided into five one-minute segments. First, the complete hybrid PINN model is trained with the given control vector $\{u_{\delta},u_{egr},u_{vgt}$}, and engine speed $n_{e}$ for time segment 0600-60 sec. The weights and biases of the network, denoted by $\bm{\theta}$, are initialized randomly. The unknown parameters for the engine, $\bm{\Lambda}:=\{A_{egrmax},A_{vgtmax},\eta_{sc},h_{tot}\},\bm{\Lambda}\in\mathbb{R}^{+}$ are initialized with a random positive value (discussed in detail in section 3). Similarly, the self-adaptive weights $\bm{\lambda}$ are also initialized randomly at the beginning of the training. Thereafter, the PINN network is trained as per the process outlined in section 2.4, and the network parameters and unknown parameters $\bm{\Lambda}$ are saved. In the TL stage, these saved parameter values are used for the initialization of network parameters (weights and biases) and unknown parameters. In Fig. 3, we illustrate and describe in detail the different stages of the proposed TL scheme.
@@
Fig. 3: Schematic for Multi-stage TL showing the network training protocol for different segments of time. In this approach, the hybrid PINN model is trained for the first segment of time ($0\sim 60$ sec). Thereafter, for the subsequent time segments ($60\sim 300$ sec in each one-minute segment), we transfer the network parameters ($\bm{\theta}$) and engine parameter estimated ($\bm{\Lambda}$) from the hybrid PINN model trained on the first time segment. Each segment uses a different control vector $\{u_{\delta},u_{egr},u_{vgt}$}, and engine speed $n_{e}$. The adaptive weights ($\bm{\lambda}$) are initialized at the start of training at epoch=0. The network weights and biases of all the layers ($\bm{\theta}$), self-adaptive weights ($\bm{\lambda}$), and unknown parameters ($\bm{\Lambda}$) are updated in Stage I training phase. In Stage II, only weights and biases of the last layer of the PINN model ($\bm{\theta_{l}}$) are trained while freezing all the hidden layers ($\bm{\theta_{h}}$). Additionally, $\bm{\lambda}$ and $\bm{\Lambda}$ continue to be updated in this phase. In Stage III, we continue to update $\bm{\theta_{l}}$ and $\bm{\Lambda}$ while the self-adaptive weights $\bm{\lambda}$ are fixed to the values obtained at the end of Stage II. In total, for predicting across a 5-minute time window, we train five models, with the model trained for segment 1 (0-60 sec) serving as the reference model from which the remaining models acquire knowledge during the TL phase.
@@ 2.6 Few-shot transfer learning
In the previous sections, first, we developed a hybrid DeepONet and PINN model and proposed a multistage TL for the inverse problem of engine parameter estimation discussed in section 1.2. In this proposed method, the model needs to be trained for the first time segment starting with random initialization of the weights and biases. To calculate the physics loss, the derivative for the output of the neural network with respect to the input needs to be calculated at each epoch. Although TL improves the computation time, the computational time required by the proposed framework is still high for field applications. One approach to improve the computational efficiency of this model is by reducing (or eliminating) the gradient calculation involved in PINN training. This may be achieved by considering the derivative to be a linear combination of some base vector. For example, in function connection [45], multi-head network [46] or reservoir computing [47], we generally do not train the hidden layer and approximate the derivative of the output as a linear combination of the derivative of the hidden layer. Desai et al. [46] incorporated one-shot inference on new tasks by training the PINN network on a training task set and then freezing all except the last layer that is trained for the new task. This approach enables one to obtain the network output as a linear combination of the output of the frozen layer with the last layer. This reduces the need for automatic differentiation to evaluate the gradients of outputs in each epoch of training with respect to inputs, which is computationally expensive. The derivative is calculated as the linear combination of the derivative of the hidden layers. We adopt this TL paradigm here, called few-shot transfer learning.
@@
Fig. 4 illustrates a schematic diagram of a few-shot TL combined with pre-trained DNN and DeepONet. Here, we note that the objectives of the pre-trained DNN and DeepONet are the same as discussed in the previous section. Hybrid PINN, using few-shot TL, may be divided into two phases. The first phase (Phase I) is an offline training phase with multiple heads for unknown variables. In this phase, we train the network offline with multiple heads, each head approximating one minute (corresponding to a different time zone) of the same output variable. The input to the network is time 0600-60 sec, normalized between [1,1][-1,1]. The first phase of offline training aims to obtain good universal basis functions (hidden layers), which can be combined to obtain appropriate output in Phase II for solving the inverse problem using PINN. In the first phase, the network is trained in a data-driven setting in an offline mode. Once the network is trained, the hidden layers (universal basis function) are transferred to the hybrid PINN model to estimate the variables in Phase II.
@@
In Phase II, we transfer the hidden layer of the multi-head network trained in the first phase. We randomly initialize the last layer ($\bm{W}_{out},b_{out}$) and set up the hybrid PINN inverse problem for one-minute duration independently. In this phase, only the last layer of the network and the unknown parameters of the equation and the self-adaptive weights are optimized. The estimated output is obtained as
@@
where $\bm{h}(t)$ is the output of the last hidden layer, $\bm{W}_{out}$ and $b_{out}$ are the weights and biases of the output layer. The derivative of the output with respect to the input variable is
@@
Since in Phase II, the hidden layers are fixed (not trained), the derivative $d\bm{h}(t)/dt$ does not change and can be kept outside the training loop for efficiency. Thus, the derivatives need to be calculated only once outside the training loop and we approximate the derivative of the output as a linear combination of the derivative of the hidden layer with the same combination matrix ($\bm{W}_{out}$). In this case, we trained only the last layer, such that we can keep the derivative outside the training loop using Eq. 20, thus reducing the computational cost. The loss function for few shot TL is the same as that of hybrid PINN given by Eq. (18). Note here that we utilize a non-linear activation function in the output layer and hence we employ the chain rule for calculating the derivative of output with respect to input. We discuss this in detail in the numerical result section 4.
Fig. 4: Schematic diagram for the few-shot TL architecture with multi-head network. The left side of the plot shows Phase I of the few-shot TL approach. In this phase, we train the multi-head networks using laboratory data where each head approximates one minute of response (corresponding to a different time zone). Input to this network is time 0600-60 sec, normalized between [1,1][-1,1]. Phase I training is a data-driven offline training. On the right side of the plot, Phase II, which involves TL, is shown. In this phase, we transfer the hidden layer and derivative of the output of the last hidden layer from the network trained in Phase I. We randomly initialize the weights and biases of the last layer along with the unknown parameters. The output and the derivative of the output are predicted using Eqs. (19) and (20), respectively. We formulate the hybrid PINN model by combining the PINN network with DeepONet and empirical DNN. The total loss ($\mathcal{L}_{\bm{\theta}}$) is calculated as the weighted sum of the physics and data loss. The weights $\lambda_{1}$ and $\lambda_{2}$ may be constant or adaptive depending on the problem setup and solution method. The trainable parameters are the weights and bias of the output layer of the PINN network ($\bm{\theta}_{out}=\{\bm{W}_{out},b_{out}\}$), the unknown parameters of the equation, the self-adaptive (in case of self-adaptive) weights. The weight and biases of the hidden layer of the PINN network, the parameters of the pre-trained network for empirical formula, and the parameters of the DeepONets remain fixed during the optimization process.
@@ 2.7 Data generation
As discussed earlier, we utilize two types of networks: (1) pre-trained DNN for the empirical formula and DeepONets (pre-trained) for the independent states ($\tilde{u}_{egr1}$, $\tilde{u}_{egr2}$ and $\tilde{u}_{vgt}$), (2) PINN networks used for online learning. Thus, two distinct types of data is required. We label the data used for training the empirical formulae networks and DeepONets for EGR and VGT states as laboratory data. The data used for solving the inverse problem through PINNs is labeled as field data, which is used to simulate a real testing scenario. A prior study [30] provides an in-depth discussion on the laboratory data required for training each DNN for empirical formulae and other necessary details. In the absence of laboratory and/or field data, we use Simulink [34] to generate both datasets presented in Table 1. Furthermore, in the case of few-shot TL, we consider the laboratory data to train the multi-head NN to obtain the body of the NN.
We consider four sets of input data (control vector and speed) to generate lab and field data and generate data for different ambient temperature and ambient pressure as shown in Table 5. The simulated data are generated with $dt=0.2$ sec (sampling frequency 5Hz5\,Hz) for training and testing of DNN for empirical formulae and DeepONet except training of $\eta_{tm}$. In the case of $\eta_{tm}$, these are generated at $dt=0.025$ sec. We trained the DNN for $f_{egr}(\mathcal{N}_{2}^{P}(:,\bm{\theta}))$ only with data from Case-I shown in Table 5. The data for training of multi-head DNN for few-shot TL are discussed later in the numerical results section. For further details regarding data generation, we refer readers to [30]. We consider PINN for each 60 sec time segment independently in our study. The initial conditions for all the state variables including for $x_{r}$ and $T_{1}$ are known for each time segment.
@@
(K) (Pa) dtdt
Case-I T65T-65 P0.100P-0.100 Set-I 1 sec Training pre-trained DNN for empirical formulae
Case-II T65T-65 $P+0.211$ Set-I 1 sec
Case-III T28T-28 P0.100P-0.100 Set-I 1 sec
@@
Case-I T65T-65 P0.100P-0.100 Set-I 1 sec Training $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$
@@
Case-I T65T-65 P0.100P-0.100 Set-I 1 sec Training $\tilde{u}_{vgt}$
@@
3 Results: Multi-stage Transfer Learning